OpenStudy (anonymous):

Write the equation for the horizontal line that contains point G(-8,8). I'm not asking for the answer. I'm just not sure exactly what the question is asking me to do.

OpenStudy (anonymous):

Equation of a line = y-y1 = m(x-x1) where m is the slope of the line. (x1,y1) = (-8,8) do you understand from here?

OpenStudy (anonymous):

Yep I already got. I wasn't sure what 'equation' they were asking for. Thank you.

OpenStudy (anonymous):

* I already got it